# Image-Based Virtual Try-On: A Survey

### v2

@@ Abstract
Image-based virtual try-on aims to synthesize a naturally dressed person image with a clothing image, which revolutionizes online shopping and inspires related topics within image generation, showing both research significance and commercial potential. However, there is a big gap between current research progress and commercial applications and an absence of comprehensive overview of this field to accelerate the development. In this survey, we provide a comprehensive analysis of the state-of-the-art techniques and methodologies in aspects of pipeline architecture, person representation and key modules such as try-on indication, clothing warping and try-on stage. We propose a new semantic criteria with CLIP, and evaluate representative methods with uniformly implemented evaluation metrics on the same dataset. In addition to quantitative and qualitative evaluation of current open-source methods, we also utilize ControlNet to fine-tune a recent large image generation model (PBE) to show future potential of large-scale models on image-based virtual try-on task. Finally, unresolved issues are highlighted and future research directions are prospected to identify key trends and inspire further exploration. The uniformly implemented evaluation metrics, dataset and collected methods will be made public available at https://github.com/little-misfit/Survey-Of-Virtual-Try-On.
Index Terms:
Virtual try-on, image synthesis, image warping, person representation, survey, AIGC.
@@ 1 Introduction
Image-based virtual Try-On is a popular research topic in the field of AI-generated content (AIGC), specifically in the domain of conditional person image generation. It enables editing, replacement, and design of clothing image content, making it highly applicable in various domains such as e-commerce platforms and short video platforms [1]. In particular, online shoppers can benefit from virtual try-on by obtaining try-on effect images of clothing, thereby enhancing their shopping experience and increasing the likelihood of successful transactions. In addition, AI Fashion has also emerged on short video platforms, where users can edit the clothes worn by characters in images or videos according to their own creativity. This allows users to explore their sense of fashion and produce a wide range of engaging images and videos.
Figure 1: A concise timeline of image-based virtual try-on milestones. Different colors indicate the main characteristic of method. Please refer to Table I for detailed comparisons.
@@
To overcome the above difficulties, tremendous efforts have been made and Fig. 1 show some representative methods on a timeline. In 2017, CAGAN[19] gave the first try by employing CycleGAN[20] to overcome the lack of training triplet data, i.e., (original person image, in-shop clothing image, try-on image), but the generation quality is far from satisfactory. Subsequently, VITON[21] creatively proposed clothing-agnostic person representation by human parsing to make up the lack of supervised training data. They constructed the basic network framework of “Try-On Indication + Cloth Warping + Try-on”, laying the foundation for further improvement on generation quality in subsequent works[22, 23, 24, 25, 26, 27, 28, 29, 30, 31]. Apart from continuous quality improvements, some new goals are desired. Trying on multiple garments attracted attention in 2019 [32], but such works are scarce due to the complex interaction between multiple garments and the lack of specialized datasets [33, 34, 35]. In order to remove the heavy reliance on human parsing, several teacher-student networks were designed [36, 37, 38] to achieve parser-free try-on at inference time. Recently, with the advancement of computational capabilities, high-resolution virtual try-on tasks has become possible, e.g., VITON-HD[39] and HR-VTON[31]. Inspired by StyleGAN[40, 41, 42, 43] and Diffusion model[44, 45, 46, 47] in the field of image generation, single-stage networks [48, 49] as well as Diffusion architecture [50] emerged for this task.
@@
  • Open challenges and future directions. According to the experimental results, we reveal a number of unresolved issues and draw important future research directions. Additionally, we utilize ControlNet [54] to fine-tune a recent large model (PBE [55]) for image-based virtual try-on, which shows the potentials and issues of popular large models on this task. We hope this review could spur the development of novel ideas towards image-based virtual try-on and its applications in industry.
The rest of this survey is organized as follows. Sec. 2 firstly gives the problem definition, and then comprehensively review the literature from multiple perspectives. Sec. 3 introduces datasets and evaluation criterion. Experimental results and analyses are presented in Sec. 4. Finally, we reveal unresolved issues Sec. 5 and prospect future directions in Sec. 6.
@@ 2.1 Overview
TABLE I: Key characteristics of representative image-based virtual try-On methods. Please refer the number of pipeline to Fig. 2 and the number of person representation is shown in Fig. 4. The three loss items are the respective losses for modules of Try-On Indication, Cloth Warping and Try-on. “-” means this item is not available or not indicated in the published work.
Model Source Pipeline
Person
Rep.
Try-On
Indication
Loss
Cloth
Warping
Loss Try-On Loss Train/Test Set
2017
CAGAN[19]
ICCVW
\@slowromancapi@
1
Cycle-GAN
Mask Combine
Mask L1
Cycle Loss
Adversarial Loss
Zalando SE
2018
SwapNet[64]
ECCV
\@slowromancapii@
2,10
DualPath U-Net
Adversarial Loss
Cross Entropy Loss
U-Net
Mask L1
Perceptual Loss
Adversarial Loss
DeepFashion
/ VITON
VITON[21]
CVPR
\@slowromancapv@
2,5,6
U-Net
Mask L1
Perceptual Loss
TPS
Mask Combine
Mask L1
Perceptual Loss
Mask TV Norm
VITON
CP-VTON[23]
ECCV
\@slowromancapvi@
2,5,6
U-Net
TPS
L1 Loss
Mask Combine
Mask L1
L1 Loss
Perceptual Loss
VITON
2019
RCW[27]
ICCVW
Use PatchGAN method to train GMM module in CP-VTON
PVT[28]
ICCVW
Add semantic information assisted training GMM in CP-VTON
LA-VTON[29]
CVPR
\@slowromancapvi@
2,5,7
U-Net
Mask L2
TPS
L1 Loss
Adversarial Loss
Grid Interval Consistency
Mask Combine
Refine Net
L1 Loss
Perceptual Loss
Adversarial Loss
VITON
UVTON[32]
ICCVW
\@slowromancapiii@
2,9
Multi-GAN
L1 Loss
Perceptual Loss
Adversarial Loss
U-Net
L1 Loss
Perceptual Loss
Adversarial Loss
VITON
VTNFP[30]
ICCV
\@slowromancapvi@
2,4,5,11
CNN
Adversarial Loss
Focal Loss[65]
TPS
L1 Loss
Adversarial Loss
Attention U-Net
Mask Combine
Mask L1
L1 Loss
Perceptual Loss
VITON
Clothflow[66]
ICCV
\@slowromancapv@
1,5,10
U-Net
Cross Entropy Loss
FlowNet
ROI Perceptual Loss
Mask L1
TV Norm
U-Net
Perceptual Loss
Style Loss
DeepFashion
/ VITON
2020
SP-VITON[25]
MTA
\@slowromancapv@
2,5,8
U-Net
Mask L1
Perceptual Loss
TPS
U-Net
Mask Combine
Perceptual Loss
Mask L1
Mask TV Norm
VITON
CP-VTON+[22]
CVPRW
\@slowromancapvi@
2,5,7
U-Net
Mask L1
TPS
Mask L1
Grid Regularization
U-Net
Mask Combine
Mask L1
L1 Loss
Perceptual Loss
VITON
WUTON[36]
ECCV
\@slowromancapiv@
3
TPS
L1 Loss
Siamese U-Net
Adversarial
L1 Loss
Perceptual Loss
VITON
ACGPN[26]
CVPR
\@slowromancapv@
2,4,11
Multi-GAN
Adversarial Loss
Cross Entropy Loss
TPS
L1 Loss
Second-order-diffe-
rence Constraint[67]
Body Part-
Composition GAN
Adversarial Loss
VITON
O-VITON[33]
CVPR
\@slowromancapv@
9,10
Auto-Encoder
Adversarial Loss
Cross Entropy Loss
Broadcast Decoder
Mask L1
L1 Loss
Perceptual Loss
VITON
2021
LM-VTON[24]
AAAI
\@slowromancapvi@
2,5,11
U-Net
Adversarial Loss
Cross Entropy Loss
TPS
L1 Loss
LandMark Loss
Perceptual Loss
Mask L1
U-Net
L1 Loss
Adversarial Loss
Perceptual Loss
VITON
MPV
OVNet[34]
CVPR
\@slowromancapv@
5,12
U-Net
Cross Entropy Loss
LSGAN Loss[68]
Multi-STN
Cascade Loss
U-Net
L1 Loss
Hinge-
Adversarial Loss[69]
Perceptual Loss
Self Made
/ VITON
CIT[70]
CVPR
\@slowromancapvi@
4,5,7
Transformer-
U-Net
Mask L1
Transformer-
TPS
L1 Loss
Grid Regularization
U-Net
Mask Combine
L1 Loss
Perceptual Loss
VITON
DCTON[71]
CVPR
\@slowromancapiv@
1,8
TPS
L1 Loss
Homography Matrix Reg
CNN
L1 Loss
Adversarial Loss
Content Preserving Loss
Perceptual Loss
VITON
VITON-HD
PFAFN[37]
CVPR
\@slowromancapiv@
5,8,12
FlowNet
Second-order
-smooth Constraint
U-Net
L1 Loss
Perceptual Loss
MPV
VITON
VITON-HD
VITON-HD[39]
CVPR
\@slowromancapv@
3,5,12
U-Net
Adversarial Loss
Cross Entropy Loss
TPS
Second-order-diffe-
rence Constraint[67]
ALIAS Generator
Adversarial Loss
Feature Matching loss
Perceptual Loss
VITON
DiOr[35]
ICCV
\@slowromancapiv@
1,5
FlowNet
Broadcast Decoder
L1 Loss
Perceptual Loss
Style Loss
Adversarial Loss
Cross Entropy Loss
Geometric Loss
DeepFashion
ZFlow[72]
ICCV
\@slowromancapvii@
4,5,7,8,9
U-Net
Cross Entropy Loss
FlowNet
L1 Loss
Perceptual Loss
Mask L1
TV Norm
U-Net
Mask Combine
L1 Loss
Perceptual Loss
Edge Loss
Segment-
Cross Entroy Loss
Densepose-
Cross Entroy Loss
UV L1
VITON
TryOnGAN[48]
ACMTG
\@slowromancapii@
1
Implicit
Transform
StyleGAN2
Editing-localization Loss
Identity Loss
Projection Loss
Garment Perceptual Loss
Self-Made
2022
Flow-Style-
VTON[38]
CVPR
\@slowromancapiv@
5,8,12
FlowNet
U-Net
L1 Loss
Garment L1 Loss
Flow-smooth Reg
Distillation Loss
VITON
DGP[49]
CVPR
\@slowromancapii@
13
Implicit
Transform
Decoder
L2 Loss
Perceptual Loss
Adversarial Loss
Attribute Loss
ESF
/ CMI MPV
HR-VITON[31]
ECCV
\@slowromancapvii@
3,8,12
Dualpath-
U-Net
Cross Entropy Loss
FlowNet
Mask L1
Adversarial Loss
Perceptual Loss
TV Norm
Mask Combine
Decoder
Adversarial Loss
Perceptual Loss
Feature Matching loss
VITON-HD
C-VTON[73]
WACV
\@slowromancapiv@
3,8
TPS
Mask L1
L1 Loss
Perceptual Loss
U-Net
Perceptual Loss
DensePose-
Adversarial Loss
Garment-
Adversarial Loss
Body-
Adversarial Loss
MPV
VITON
SDAFN[74]
ECCV
\@slowromancapiv@
3,4
Multi-
FlowNet
Decoder
Multi Layer:
L1 Loss
Style Loss
Perceptual Loss
MPV
VITON
Model source Pipeline
Person
Rep.
Try-On
Indication
Loss
Cloth
Warping
Loss Try-On Loss Train/Test Set
2022
3D-GCL[75]
NeurIPS
\@slowromancapiv@
9
FlowNet
L1 Loss
Perceptual Loss
3D Regularization
StyleGAN2
L1 Loss
Perceptual Loss
Adversarial Loss
MPV
DeepFashion
2023
GP-VTON[76]
CVPR
\@slowromancapvii@
5,8,12
Dualpath-
U-Net
Cross Entropy Loss
Adversarial Loss
Multi-
FlowNet
L1 Loss
Mask L1
Perceptual Loss
Second-order
-smooth Constraint
Mask Combine
+U-Net
L1 Loss
Mask L1
Adversarial Loss
Perceptual Loss
VITON-HD
Dress Code
SAL-VTON[77]
CVPR
\@slowromancapvii@
1,13
Multi-
FlowNet
L1 Loss
LandMark Loss
Perceptual Loss
Smoothness Norm[78]
U-Net
Adversarial Loss
Perceptual Loss
Feature Matching loss
VITON
VITON-HD
TryOn
Diffusion[50]
CVPR
\@slowromancapi@
4
Large
Model
Diffusion Denosing
Loss
MPV
VITON
It is worth noting that the above three steps are not necessarily present at the same time, and there is no strict order. Table I summarizes representative methods and we will discuss the key designs in the following subsections.
Figure 2: Basic pipelines of image-based virtual try-on. Pipelines \@slowromancapi@ and \@slowromancapii@ are both single-stage approaches, where pipeline \@slowromancapi@ utilizes a single generator to directly generate the try-on image, while pipeline \@slowromancapii@ aligns features in the feature domain before generating the try-on image. Pipelines \@slowromancapiii@ and \@slowromancapiv@ are both two-stage pipelines, where the former utilizes person representation as the bridge while the later uses warped clothing. Pipelines \@slowromancapv@ and \@slowromancapvi@ are three-stage pipelines, which differ in the order of Try-On Indication and Cloth Warping. Pipeline \@slowromancapvii@ is an improvement over \@slowromancapv@ and \@slowromancapvi@, which simultaneously performs Try-On Indication and Cloth Warping, with/without both modules communicating during the data processing. For detailed explanations of the Try-On Indication, Cloth Warping, and Try-On module, please refer to Section 2.4, 2.5, 2.6.
(a) Teacher-Student I
(b) Teacher-Student II
(c) Cycle-GAN Structure
@@ 2.2 Pipelines
In the virtual try-on pipeline, the selection and placement of the aforementioned three modules have a significant impact on the final try-on results. As shown in Fig. 2, the basic pipeline structure can be categorized into seven types. Type \@slowromancapi@ and \@slowromancapii@ are one-stage pipelines and the later one additionally introduces feature alignment. Type \@slowromancapiii@ and \@slowromancapiv@ are two-stage pipelines which respectively utilize person representation and warped clothing as intermediate generation for further optimization. The rest types are three-stage pipelines, where type \@slowromancapv@ and \@slowromancapvi@ differ in the order of Try-On Indication and Cloth Warping modules while type \@slowromancapvii@ simultaneously optimizes these two modules. The pipeline choice of representative methods can be found in Table I, with no obvious preference in the development trend.
Supplemental to the basic try-on image generation pipeline, Fig. 3 shows other structures such as Teacher-Student network [36, 38, 37] and Cycle-GAN [20, 77]. The Teacher-Student architecture is mainly designed for training parser-free try-on network and Fig. 3a shows the straightforward implementation [36]. PFAFN and Style-Flow-VTON [38, 37] further improve it to Fig. 3b where the synthesized image generated by the teacher network $I_{try-on}$ is used as the input of the student network. Compared with type 1, type 2 provides more reliable supervision with the ground truth in case that the teacher network generates poor results. Similarly, the adoption of Cycle-GAN [20, 77] (Fig. 3c) also shows a strategy for using cycle consistency to enhance the supervision in try-on.

### v3

@@ Abstract
Image-based virtual try-on aims to synthesize a naturally dressed person image with a clothing image, which revolutionizes online shopping and inspires related topics within image generation, showing both research significance and commercial potential. However, there is a gap between current research progress and commercial applications and an absence of comprehensive overview of this field to accelerate the development. In this survey, we provide a comprehensive analysis of the state-of-the-art techniques and methodologies in aspects of pipeline architecture, person representation and key modules such as try-on indication, clothing warping and try-on stage. We propose a new semantic criteria with CLIP, and evaluate representative methods with uniformly implemented evaluation metrics on the same dataset. In addition to quantitative and qualitative evaluation of current open-source methods, unresolved issues are highlighted and future research directions are prospected to identify key trends and inspire further exploration. The uniformly implemented evaluation metrics, dataset and collected methods will be made public available at https://github.com/little-misfit/Survey-Of-Virtual-Try-On.
keywords:
Virtual try-on, image synthesis, image warping, person representation, survey, AIGC
@@ 1 Introduction
Image-based virtual try-on is a popular research topic in the field of AI-generated content (AIGC), specifically in the domain of conditional person image generation. It enables editing, replacement, and design of clothing image content, making it highly applicable in various domains such as e-commerce platforms and short video platforms. In particular, online shoppers can benefit from virtual try-on by obtaining try-on effect images of clothing, thereby enhancing their shopping experience and increasing the likelihood of successful transactions. In addition, AI Fashion has also emerged on short video platforms, where users can edit the clothes worn by characters in images or videos according to their own creativity. This allows users to explore their sense of fashion and produce a wide range of engaging images and videos.
Figure 1: A concise timeline of image-based virtual try-on milestones. Different colors indicate the main characteristic of method. Please refer to Table 1 for detailed comparisons.
@@
To overcome the above difficulties, tremendous efforts have been made and Fig. 1 show some representative methods on a timeline. In 2017, CAGAN [50] gave the first try by employing CycleGAN [141] to overcome the lack of training triplet data, i.e., (original person image, in-shop clothing image, try-on image), but the generation quality is far from satisfactory. Subsequently, VITON [37] creatively proposed clothing-agnostic person representation by human parsing to make up the lack of supervised training data. They constructed the basic network framework of “Try-On Indication + Cloth Warping + Try-on”, laying the foundation for further improvement on generation quality in subsequent works [82, 119, 75, 111, 130, 3, 4, 60, 135, 61]. Apart from continuous quality improvements, some new goals are desired. Trying on multiple garments attracted attention in 2019 [58], but such works are scarce due to the complex interaction between multiple garments and the lack of specialized datasets [85, 65, 22]. In order to remove the heavy reliance on human parsing, several teacher-student networks were designed [47, 31, 40] to achieve parser-free try-on at inference time. Recently, with the advancement of computational capabilities, high-resolution virtual try-on tasks has become possible, e.g., VITON-HD [18] and HR-VTON [61]. Inspired by StyleGAN [54, 55, 73] and Diffusion model [43, 87, 112, 70] in the field of image generation, single-stage networks [62, 27] as well as Diffusion architecture [142, 56, 136] emerged for this task.
@@
  • Open challenges and future directions. According to the experimental results, we reveal a number of unresolved issues and draw important future research directions. We hope this review could spur the development of novel ideas towards image-based virtual try-on and its applications in industry.
The rest of this survey is organized as follows. Sec. 2 firstly gives the problem definition, and then comprehensively review the literature from multiple perspectives. Sec. 3 introduces datasets and evaluation criterion. Experimental results and analyses are presented in Sec. 4. Finally, we reveal unresolved issues in Sec. 5 and prospect future directions in Sec. 6.
Table 1: Key characteristics of representative image-based virtual try-On methods. Please refer the number of pipeline to Fig. 2 and the number of person representation is shown in Fig. 4. The three loss items are the respective losses for modules of Try-On Indication, Cloth Warping and Try-on. “-” means this item is not available or not indicated in the published work.
Year Model Source Pipeline Person Try-On Loss Cloth Loss Try-On Loss Train/Test
Rep. Indication Per L1 Adv Ent Warping Per L1 Adv Reg Per L1 Adv Oths Dataset
2017
CAGAN
[50]
ICCVW \@slowromancapi@
1
Cycle-GAN
Mask Combine
Zalando SE
2018
VITON
[37]
CVPR \@slowromancapv@
2,5,6
U-Net TPS - Mask Combine VITON
CP-VTON
[119]
ECCV \@slowromancapvi@
2,5,6
U-Net - TPS Mask Combine VITON
SwapNet
[96]
ECCV \@slowromancapii@
2,10
Dualpath-
U-Net
- - U-Net
DeepFashion
/ VITON
2019
LA-VTON
[60]
CVPR \@slowromancapvi@
2,5,7
U-Net TPS
Mask Combine
Refine Net
VITON
VTNFP
[135]
ICCV \@slowromancapvi@
2,4,5,11
CNN TPS
Attention U-Net
Mask Combine
VITON
Clothflow
[38]
ICCV \@slowromancapv@
1,5,10
U-Net FlowNet
U-Net
DeepFashion
/ VITON
UVTON
[58]
ICCVW \@slowromancapiii@
2,9
Multi-GAN - - U-Net VITON
2020
SP-VITON
[111]
MTAP \@slowromancapv@
2,5,8
U-Net TPS -
U-Net
Mask Combine
VITON
ACGPN
[130]
CVPR \@slowromancapv@
2,4,11
Multi-GAN TPS
Body Part-
Composition GAN
VITON
O-VITON
[85]
CVPR \@slowromancapv@
9,10
Auto-Encoder - -
Broadcast Decoder
VITON
CP-VTON+
[82]
CVPRW \@slowromancapvi@
2,5,7
U-Net TPS
U-Net
Mask Combine
VITON
WUTON
[47]
ECCV \@slowromancapiv@
3
- - TPS
Siamese U-Net
VITON
2021
TryOnGAN
[62]
TOG \@slowromancapii@
1
Implicit
Transform
StyleGAN2
Self-Made
LM-VTON
[75]
AAAI \@slowromancapvi@
2,5,11
U-Net TPS
U-Net
VITON
MPV
OVNet
[65]
CVPR \@slowromancapv@
5,12
U-Net Multi-STN
U-Net
Self Made
/ VITON
CIT
[97]
CVPR \@slowromancapvi@
4,5,7
Transformer-
U-Net
Transformer-
TPS
U-Net
Mask Combine
VITON
DCTON
[30]
CVPR \@slowromancapiv@
1,8
- - TPS
CNN
VITON
PFAFN
[31]
CVPR \@slowromancapiv@
5,8,12
- - FlowNet
U-Net
MPV
VITON
VITON-HD
[18]
CVPR \@slowromancapv@
3,5,12
U-Net TPS
ALIAS Generator
VITON-HD
DiOr
[22]
ICCV \@slowromancapiv@
1,5
- - FlowNet -
Broadcast Decoder
DeepFashion
ZFlow
[20]
ICCV \@slowromancapvii@
4,5,7,8,9
U-Net FlowNet
U-Net
Mask Combine
VITON
2022
C-VTON
[26]
WACV \@slowromancapiv@
3,8
- - TPS
U-Net
MPV
VITON
Flow-Style-VTON
[40]
CVPR \@slowromancapiv@
5,8,12
- - FlowNet -
U-Net
VITON
DGP
[27]
CVPR \@slowromancapii@
13
Implicit
Transform
Decoder
ESF
/ CMI MPV
RT-VTON
[131]
CVPR \@slowromancapv@
1,5,10
CNN
TPS
U-Net
VITON
HR-VITON
[61]
ECCV \@slowromancapvii@
3,8,12
Dualpath-
U-Net
FlowNet
Mask Combine
Decoder
VITON-HD
SDAFN
[5]
ECCV \@slowromancapiv@
3,4
Multi-
FlowNet
Decoder
MPV
VITON
3D-GCL
[46]
NeurIPS \@slowromancapiv@
9
- - FlowNet
StyleGAN2
MPV
DeepFashion
2023
POVNet
[66]
TPAMI \@slowromancapv@
2,5,12
U-Net Multi-STN U-Net
Dress-Code
VITON-HD
GP-VTON
[126]
CVPR \@slowromancapvii@
5,8,12
Dualpath-
U-Net
Multi-
FlowNet
Mask Combine
+U-Net
VITON-HD
Dress Code
SAL-VTON
[128]
CVPR \@slowromancapiv@
1,13
Multi-
FlowNet
U-Net
VITON
VITON-HD
TryOnDiffusion
[142]
CVPR \@slowromancapi@
4
- - - - Diffusion
MPV
VITON
KGI
[69]
ICCV \@slowromancapi@
3,5,12,13
Auto-Encoder TPS Diffusion VITON-HD
COTTON
[15]
ICCV \@slowromancapi@
4,5,10
U-Net STN U-Net
COTTON
VITON-HD
Ladi-VTON
[84]
ACM MM \@slowromancapi@
3,5
- - - - Diffusion
Dress-Code
VITON-HD
DCI-VTON
[34]
ACM MM \@slowromancapiv@
3, 8,12
- - FlowNet Diffusion
Dress-Code
VITON-HD
2024
SD-VITON
[107]
AAAI \@slowromancapvii@
3,8,12
Dualpath-
U-Net
FlowNet
Mask Combine
Decoder
VITON-HD
Stable-VITON
[56]
CVPR \@slowromancapi@
3,8
Diffusion
Dress-Code
VITON-HD
CAT-DM
[136]
CVPR \@slowromancapi@
3,8
- - - - Diffusion
Dress-Code
VITON-HD
TPD
[133]
CVPR \@slowromancapiii@
3,5,8
Diffusion - - - Diffusion
VITON
VITON-HD
Figure 2: Basic pipelines of image-based virtual try-on. Pipelines \@slowromancapi@ and \@slowromancapii@ are both single-stage approaches, where pipeline \@slowromancapi@ utilizes a single generator to directly generate the try-on image, while pipeline \@slowromancapii@ aligns features in the feature domain before generating the try-on image. Pipelines \@slowromancapiii@ and \@slowromancapiv@ are both two-stage pipelines, where the former utilizes person representation as the bridge while the later uses warped clothing. Pipelines \@slowromancapv@ and \@slowromancapvi@ are three-stage pipelines, which differ in the order of Try-On Indication and Cloth Warping. Pipeline \@slowromancapvii@ is an improvement over \@slowromancapv@ and \@slowromancapvi@, which simultaneously performs Try-On Indication and Cloth Warping.
@@ 2.1 Overview
It is worth noting that the above three steps are not necessarily present at the same time, and there is no strict order. Table 1 summarizes representative methods and we will discuss the key designs in the following subsections. The statistics of image-based virtual try-on methods are summarized from the literature listed in Table 1 and the following methods [3, 4, 2, 45, 64, 39, 103, 98, 28, 80, 93, 104, 49, 95, 44, 113, 24, 19, 88, 14, 127, 76, 140, 53, 59, 91, 125, 123, 63, 72, 121, 124, 29, 6, 92, 134, 139, 71, 86, 90, 106, 1].
@@ 2.2 Pipelines
In the virtual try-on pipeline, the selection and placement of the aforementioned three modules have a significant impact on the final try-on results. As shown in Fig. 2, the basic pipeline structure can be categorized into seven types. Type \@slowromancapi@ and \@slowromancapii@ are one-stage pipelines and the later one additionally introduces feature alignment. Type \@slowromancapiii@ and \@slowromancapiv@ are two-stage pipelines which respectively utilize person representation and warped clothing as intermediate generation for further optimization. The rest types are three-stage pipelines, where type \@slowromancapv@ and \@slowromancapvi@ differ in the order of Try-On Indication and Cloth Warping modules while type \@slowromancapvii@ simultaneously optimizes these two modules. The pipeline choice of representative methods can be found in Table 1, with no obvious preference in the development trend.
Supplemental to the basic try-on image generation pipeline, Fig. 3 shows other structures such as Teacher-Student network [47, 40, 31] and Cycle-GAN [141, 128]. The Teacher-Student architecture is mainly designed for training parser-free try-on network and Fig. 3(a) shows the straightforward implementation [47]. PFAFN and Style-Flow-VTON [40, 31] further improve it to Fig. 3(b) where the synthesized image generated by the teacher network $I_{try-on}$ is used as the input of the student network. Compared with type \@slowromancapi@, type \@slowromancapii@ provides more reliable supervision with the ground truth in case that the teacher network generates poor results. Similarly, the adoption of Cycle-GAN [141, 128] (Fig. 3(c)) also shows a strategy for using cycle consistency to enhance the supervision in try-on.